OpenStudy (anonymous):

Solution to x+3y=3 & 2x-y=-3

OpenStudy (anonymous):

The easiest way I can think of is to single out the y variable. Then graph the two equations. x+3y=3 You can subtract x from both sides. Which leaves you with: 3y=3-x Then divide both sides by 3, so that y is fully alone. Then do the same thing with the second equation. Afterwards, you'll be able to put them into any graphing calculator and find the solutions.

OpenStudy (anonymous):

I would solve for one of the variables then plug it into the other equation x=-3y+3 and 2x-y=-3 so it would be 2(-3y+3)-y=-3 making it so -6y+6-y=-3 than solve for y once that is done plug y in and solve for x doing the same thing

OpenStudy (anonymous):

That could also work. It just depends on which on you are more comfortable with.